# Spectral Masking and Interpolation Attack (SMIA): A Black-box Adversarial Attack against Voice Authentication and Anti-Spoofing Systems

### v3

@@ 1. Introduction
Voice authentication has emerged as a widely adopted biometric technology due to its ease of use and flexibility in enhancing security within various scenarios, including mobile devices, virtual assistants, secure communication systems, financial service institutions, and law enforcement and jurisdictional departments (singh2012applications; algabri2017automatic; rose2006technical). These systems analyze an individual’s voice to create a distinct vocal signature used for identity verification.
@@
Compared to passwords or access cards, voice authentication offers reliable user verification, reduces fraud risk, and simplifies the user experience by eliminating the need to remember multiple credentials and avoiding password-related vulnerabilities. It is also more stable and cost-effective than other biometric systems. Unlike facial recognition, which requires frequent updates due to changes in appearance, voice characteristics remain relatively consistent after adulthood. Additionally, it relies on affordable microphones and requires less storage than high-resolution cameras and facial data. Additionally, it is easier to integrate with cloud services and multiple devices (renz2023authentication; springer2023twofactor; techlocity2021biometrics; khan2010comparative).
@@
Recent advancements in deep learning have revolutionized the Voice Authentication System (VAS), significantly improving accuracy, robustness, and scalability. Modern deep learning-based approaches, such as Deep Neural Network (DNN) models, have enhanced the field by creating digital voice fingerprints that serve as reference points for authentication, for example, Deep Speaker (deepspeaker) and X-Vectors (xvectors). Despite these improvements, VAS remain highly vulnerable to various forms of adversarial attacks (lan2022adversarial).
@@
In parallel, deep learning has transformed the field of synthetic speech generation, making it easier than ever to replicate someone’s voice with near-human indistinguishability. Technologies like text-to-speech (TTS) (fish-speech-v1.4; qin2023openvoice; amazon_polly) and voice conversion (VC) (li2023freevc; huang2024diffvc+) can generate highly realistic speech using minimal input data. These tools, while impressive, have raised significant concerns regarding their misuse in security-sensitive applications, particularly when synthetic speech is used to deceive or bypass VAS (shonesy2015voice; alali2025partial).
@@
To combat these deepfake attacks, various countermeasures (CMs), like RawNet2 (rawnet2), RawGAT-ST (rawgat), and RawPC-Darts (rawdarts), have been developed to distinguish between genuine and synthetic speech. However, attackers have engineered a new class of threat, known as adversarial spoofing attacks, designed specifically to deceive these CMs. This hybrid approach begins by generating a convincing spoofed audio sample using TTS or VC. Subsequently, a carefully crafted and often imperceptible adversarial perturbation is added to this synthetic speech. The resulting signal has a dual objective: bypassing the primary VAS by mimicking the target speaker while simultaneously fooling the CM into classifying the spoofed audio as authentic, posing a compounded threat to security protocols (hai2023sifdetectcracker; kassis2023breaking).
@@
A compelling example is SiFDetectCracker (hai2023sifdetectcracker), an adversarial attack that can deceive SOTA CMs by only modifying speaker-irrelevant features, such as background noise and silent segments. This method achieved an average success rate of over 80% without altering the speaker’s actual voice, demonstrating that detectors often over-rely on secondary acoustic cues. However, their impact on VAS or hybrid systems, which consist of VAS and CMs, remains unexplored. This highlights a critical gap: the need for robust attacks that are effective in real-world scenarios.
@@
Recent attacks on VAS have shifted focus from fooling speaker verification systems to breaking their CMs. Kassis et al. (kassis2023breaking) introduced a model-agnostic attack that bypasses these CMs without relying on query-based optimization, but by applying a series of universal signal-processing transformations to spoofed audio. Their approach, which manipulates cues like silence and spectral energy distribution, is capable of bypassing the entire VA stack. However, their results reveal a significant performance drop against more modern and robust defenses; for instance, the success rate fell to just 15.55% against a combined defense using the AWS Voice ID (amazon2025voiceid) and RawPC-Darts (rawdarts). This inconsistency highlights that while targeting universal flaws in CMs is a promising direction, a critical gap remains in developing attacks that can reliably defeat the current generation of SOTA, hardened VA systems.
@@ 2.2. Anti-Spoofing Countermeasures (CMs)
Fake voice detection has evolved considerably over the past decade. Early systems primarily relied on conventional machine learning pipelines that extracted handcrafted acoustic features such as Mel-Frequency Cepstral Coefficients (MFCCs) and Constant-Q Cepstral Coefficients (CQCCs), which were then passed into classifiers such as Gaussian Mixture Models (GMMs) and Support Vector Machines (SVMs) (patel2015svm; de2017intro). While effective at the time, these methods struggled with generalization against unseen spoofing attacks.
The field has shifted towards deep learning methods (lavrentyeva2019stc), with SOTA now driven by Self-Supervised Learning (SSL) models like Wav2Vec2.0 (baevski2020wav2vec) and WavLM (chen2022wavlm). WavLM has shown superior performance by capturing fine-grained artifacts commonly found in deepfakes. However, pre-training these large-scale SSL models takes a lot of time and computational resources. These SSL embeddings are typically followed by smaller, task-specific back-end classifiers such as the Light Convolutional Neural Network (LCNN) (lavrentyeva2017audio) or the AASIST model (jung2022aasist), which specialize in spoofing detection.
In parallel, another leading paradigm involves fully end-to-end systems operating directly on the raw waveform, removing the need for explicit feature engineering. The RawNet family exemplifies this approach: RawNet2 (rawnet2) demonstrated strong performance in the ASVspoof 2019 challenge (Todisco2019ASVspoof), outperforming its predecessor and serving as a competitive baseline. Other end-to-end architectures have emerged as SOTA alternatives. RawPC-Darts (rawdarts) applies differentiable neural architecture search (NAS) to optimize network design directly on raw audio, significantly improving efficiency and detection accuracy. Meanwhile, RawGAT-ST (rawgat) integrates graph attention networks with spectro-temporal modeling to enhance the robustness of spoof detection against unseen attacks. Both approaches highlight the shift towards more adaptive and structurally optimized architectures in the fight against audio deepfakes.
@@ 2.3. Attacks Taxonomies
Saxena et al. (shonesy2015voice) 2015 SV Black-box (Voice morphing) SV, Human listeners
Alali et al. (alali2025partial) 2025 SV Black-box (Zero-shot TTS / VC) SV, Human listeners
@@
Liu et al. (liu2019adversarial) 2019 CM White-box (FGSM/PGD) ASVspoof CM
Gomez-Alanis et al. (gomez2022adversarial) 2022 SV + CM White-box (ABTN) ASVspoof systems
Gomez-Alanis et al. (gomez2022ganba) 2022 SV + CM White-box (GANBA) ASVspoof systems
Hai et al. (hai2023sifdetectcracker) 2023 CM Black-box CM
Kassis et al. (kassis2023breaking) 2023 SV + CM Black-box (OTA spoofing) VA systems (SV + CM)
@@ 2.4. Deepfake Voice Attacks
Deepfake voice attacks involve the use of speech synthesis (TTS or VC) to mimic the vocal identity of a target speaker. Modern deepfake tools can operate in real time, increasing the feasibility of live impersonation (ren2019fastspeech; ren2020fastspeech; kim2020glow; kaneko2018cyclegan; kameoka2018stargan; qian2019autovc; li2023freevc; qin2023openvoice; wang2024maskgct). Recent advancements in zero-shot voice cloning enable attackers to generate convincing deepfakes with minimal reference audio, which lowers the barrier to entry. Currently, researchers produce CMs to prevent such attacks (almutairi2022review).
One of the earliest demonstrations of AI-powered voice spoofing was conducted by Saxena et al. (shonesy2015voice). Using a basic voice morphing tool, they were able to reconstruct a target speaker’s voice with only a few minutes of enrollment data. The spoofed voices were tested against both automated speaker verification systems and human listeners. Results revealed significant vulnerabilities: SOTA SRS failed to reject 80–90% of spoofed attempts, while human listeners correctly detected fake voices only about 50% of the time—particularly when the attacker mimicked familiar individuals.
Building on this, Alali et al. (alali2025partial) explored partial voice attacks in real-world contexts. Their study demonstrated that both humans and SRS remain vulnerable when exposed to partial speech generated through zero-shot TTS and VC models. The findings highlighted a high success rate in bypassing verification mechanisms and deceiving listeners. However, the study’s broad evaluation across different attacks and platforms lacked a deeper analysis of specific conditions, such as noisy environments, language variations, or system-level defenses, limiting its applicability for improving individual systems.
@@ 2.5.1. White-box Attacks
Early work by Liu et al. (liu2019adversarial) applied white-box adversarial techniques, notably the Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD), to CM systems. Even minimal perturbations (with $\epsilon$ below human perceptibility) caused error rates to spike dramatically, reaching 85–95% in some cases. These findings showed that, when attackers have access to model gradients, the CM’s ability to detect spoofed speech can be almost completely undermined, effectively reducing performance to random guessing.
Expanding on this, Gomez-Alanis et al. (gomez2022adversarial) introduced the Adversarial Biometrics Transformation Network (ABTN), a trainable generator that optimizes perturbations using both CM and ASV losses. Compared to FGSM, ABTN produced perturbations that generalized better across spoof types while still preserving speaker identity, thereby fooling CMs without triggering alarms in ASV systems. The method improved attack equal error rates (EER) by 27–51% (absolute) across ASVspoof 2019 logical and physical access protocols (Todisco2019ASVspoof).
In a related study, Alanis et al. (gomez2022ganba) leveraged a generative adversarial network (GAN) to deceive both CM and speaker verification (SV) systems while maintaining speaker identity. Although highly effective, this approach required access to the target or a surrogate model and concentrated solely on attack generation, without proposing defensive strategies—limiting its practical applicability.
@@ 2.5.2. Black-box Attacks
In the domain of black-box adversarial attacks, Hai et al. (hai2023sifdetectcracker) proposed SiFDetectCracker, a framework targeting CMs. Their method generates perturbations for speaker-irrelevant features (SiFs), specifically the background noise and the mute parts of an audio sample. The framework was evaluated against several advanced detectors, including RawNet2(rawnet2), RawGAT-ST(rawgat), RawPC-Darts (rawdarts), and Deep4SNet (ballesteros2021deep4snet), using the ASVspoof 2019 dataset (Todisco2019ASVspoof). The attack proved highly effective, achieving an average success rate of 82.2% by showing that many detection systems are sensitive to these non-speech features. While the paper presents a novel attack vector, ablation studies showed that the attack’s performance drops significantly when either time or noise perturbations are removed, indicating a reliance on both factors. Furthermore, the transferability between models with different architectures was low, a limitation our work aims to address.
While the work by Kassis and Hengartner (kassis2023breaking) presents a novel methodology for bypassing VAS, its real-world applicability is limited by key limitations in its evaluation. Their results against combination systems rely on a cumulative metric, which is not descriptive enough to show the state or the actual effectiveness of their attack against these more complex defenses. Furthermore, the attack’s performance dropped significantly when tested against SOTA anti-spoofing systems, like RawPC-Darts (rawdarts) and RawGAT-ST (rawgat), raising doubts about its effectiveness against modern countermeasures. Finally, the designated-app attack vector relies on the significant assumption that an attacker can use a rooted phone to inject audio, a prerequisite that is often blocked by the security measures of financial applications.
@@ 3.1. Threat Model
  • Sample Generation: They can generate a spoofed audio sample using any voice-cloning tool (e.g., Fish Speech APIs (fish-speech-v1.4) or methods used in ASVspoof 2019 (Todisco2019ASVspoof)).
@@
  • Score and Label: The system returns both a continuous confidence score and the final decision (like Azure SV (microsoft_speaker)).
@@ 3.2. Overview
As shown in Figure 2, the attack pipeline starts by generating a base spoofed audio. This is accomplished by feeding a victim’s voice sample and a target prompt into a voice synthesis model. Our framework is tool-independent—the attacker can employ Fish Speech APIs (fish-speech-v1.4), ASVSpoof 2019 challenge tools (Todisco2019ASVspoof), or any modern voice cloning system.
@@ 4. Experimental Setup
Our experimental setup leverages the LibriSpeech (panayotov2015librispeech) and ASVSpoof2019 (Todisco2019ASVspoof) datasets to assess system performance. We benchmark our proposed attacks against three state-of-the-art anti-spoofing models: RawNet2 (rawnet2), RawGAT-ST(rawgat), and RawPC-Darts(rawdarts). For the voice authentication stage, we employ three distinct systems: X-Vectors (xvectors), DeepSpeaker (deepspeaker), and the commercial Azure Speaker Verification API (microsoft_speaker). All experiments were conducted on a single NVIDIA T4 GPU. The evaluation is structured to analyze performance under various configurations:

### v4

@@ 1. Introduction
Voice authentication has emerged as a widely adopted biometric technology due to its ease of use and flexibility in enhancing security within various scenarios, including mobile devices, virtual assistants, secure communication systems, financial service institutions, and law enforcement and jurisdictional departments (Singh et al., 2012; Algabri et al., 2017; Rose, 2006). These systems analyze an individual’s voice to create a distinct vocal signature used for identity verification.
@@
Compared to passwords or access cards, voice authentication offers reliable user verification, reduces fraud risk, and simplifies the user experience by eliminating the need to remember multiple credentials and avoiding password-related vulnerabilities. It is also more stable and cost-effective than other biometric systems. Unlike facial recognition, which requires frequent updates due to changes in appearance, voice characteristics remain relatively consistent after adulthood. Additionally, it relies on affordable microphones and requires less storage than high-resolution cameras and facial data. Additionally, it is easier to integrate with cloud services and multiple devices (Renz et al., 2023; Springer, 2023; Techlocity, 2021; Khan and Zahid, 2010).
@@
Recent advancements in deep learning have revolutionized the Voice Authentication System (VAS), significantly improving accuracy, robustness, and scalability. Modern deep learning-based approaches, such as Deep Neural Network (DNN) models, have enhanced the field by creating digital voice fingerprints that serve as reference points for authentication, for example, Deep Speaker (Li et al., 2017) and X-Vectors (Snyder et al., 2018). Despite these improvements, VAS remain highly vulnerable to various forms of adversarial attacks (Lan et al., 2022).
@@
In parallel, deep learning has transformed the field of synthetic speech generation, making it easier than ever to replicate someone’s voice with near-human indistinguishability. Technologies like text-to-speech (TTS) (Liao et al., 2024; Qin et al., 2023; Amazon Web Services, 2025) and voice conversion (VC) (Li et al., 2023; Huang et al., 2024) can generate highly realistic speech using minimal input data. These tools, while impressive, have raised significant concerns regarding their misuse in security-sensitive applications, particularly when synthetic speech is used to deceive or bypass VAS (Shonesy, 2015; Alali and Theodorakopoulos, 2025).
@@
To combat these deepfake attacks, various countermeasures (CMs), like RawNet2 (Tak et al., 2021b), RawGAT-ST (Tak et al., 2021a), and RawPC-Darts (Ge et al., 2021), have been developed to distinguish between genuine and synthetic speech. However, attackers have engineered a new class of threat, known as adversarial spoofing attacks, designed specifically to deceive these CMs. This hybrid approach begins by generating a convincing spoofed audio sample using TTS or VC. Subsequently, a carefully crafted and often imperceptible adversarial perturbation is added to this synthetic speech. The resulting signal has a dual objective: bypassing the primary VAS by mimicking the target speaker while simultaneously fooling the CM into classifying the spoofed audio as authentic, posing a compounded threat to security protocols (Hai et al., 2023; Kassis and Hengartner, 2023).
@@
A compelling example is SiFDetectCracker (Hai et al., 2023), an adversarial attack that can deceive SOTA CMs by only modifying speaker-irrelevant features, such as background noise and silent segments. This method achieved an average success rate of over 80% without altering the speaker’s actual voice, demonstrating that detectors often over-rely on secondary acoustic cues. However, their impact on VAS or hybrid systems, which consist of VAS and CMs, remains unexplored. This highlights a critical gap: the need for robust attacks that are effective in real-world scenarios.
@@
Recent attacks on VAS have shifted focus from fooling speaker verification systems to breaking their CMs. Kassis et al. (Kassis and Hengartner, 2023) introduced a model-agnostic attack that bypasses these CMs without relying on query-based optimization, but by applying a series of universal signal-processing transformations to spoofed audio. Their approach, which manipulates cues like silence and spectral energy distribution, is capable of bypassing the entire VA stack. However, their results reveal a significant performance drop against more modern and robust defenses; for instance, the success rate fell to just 15.55% against a combined defense using the AWS Voice ID (Amazon, 2025) and RawPC-Darts (Ge et al., 2021). This inconsistency highlights that while targeting universal flaws in CMs is a promising direction, a critical gap remains in developing attacks that can reliably defeat the current generation of SOTA, hardened VA systems.
@@ 2.2. Anti-Spoofing Countermeasures (CMs)
Fake voice detection has evolved considerably over the past decade. Early systems primarily relied on conventional machine learning pipelines that extracted handcrafted acoustic features such as Mel-Frequency Cepstral Coefficients (MFCCs) and Constant-Q Cepstral Coefficients (CQCCs), which were then passed into classifiers such as Gaussian Mixture Models (GMMs) and Support Vector Machines (SVMs) (Patel and Patil, 2015; De Leon et al., 2017). While effective at the time, these methods struggled with generalization against unseen spoofing attacks.
The field has shifted towards deep learning methods (Lavrentyeva et al., 2019), with SOTA now driven by Self-Supervised Learning (SSL) models like Wav2Vec2.0 (Baevski et al., 2020) and WavLM (Chen et al., 2022). WavLM has shown superior performance by capturing fine-grained artifacts commonly found in deepfakes. However, pre-training these large-scale SSL models takes a lot of time and computational resources. These SSL embeddings are typically followed by smaller, task-specific back-end classifiers such as the Light Convolutional Neural Network (LCNN) (Lavrentyeva et al., 2017) or the AASIST model (Jung et al., 2022), which specialize in spoofing detection.
In parallel, another leading paradigm involves fully end-to-end systems operating directly on the raw waveform, removing the need for explicit feature engineering. The RawNet family exemplifies this approach: RawNet2 (Tak et al., 2021b) demonstrated strong performance in the ASVspoof 2019 challenge (et al., 2019), outperforming its predecessor and serving as a competitive baseline. Other end-to-end architectures have emerged as SOTA alternatives. RawPC-Darts (Ge et al., 2021) applies differentiable neural architecture search (NAS) to optimize network design directly on raw audio, significantly improving efficiency and detection accuracy. Meanwhile, RawGAT-ST (Tak et al., 2021a) integrates graph attention networks with spectro-temporal modeling to enhance the robustness of spoof detection against unseen attacks. Both approaches highlight the shift towards more adaptive and structurally optimized architectures in the fight against audio deepfakes.
@@ 2.3. Attacks Taxonomies
Saxena et al. (Shonesy, 2015) 2015 SV Black-box (Voice morphing) SV, Human listeners
Alali et al. (Alali and Theodorakopoulos, 2025) 2025 SV Black-box (Zero-shot TTS / VC) SV, Human listeners
@@
Liu et al. (Liu et al., 2019) 2019 CM White-box (FGSM/PGD) ASVspoof CM
Gomez-Alanis et al. (Gomez-Alanis et al., 2022a) 2022 SV + CM White-box (ABTN) ASVspoof systems
Gomez-Alanis et al. (Gomez-Alanis et al., 2022b) 2022 SV + CM White-box (GANBA) ASVspoof systems
Hai et al. (Hai et al., 2023) 2023 CM Black-box CM
Kassis et al. (Kassis and Hengartner, 2023) 2023 SV + CM Black-box (OTA spoofing) VA systems (SV + CM)
@@ 2.4. Deepfake Voice Attacks
Deepfake voice attacks involve the use of speech synthesis (TTS or VC) to mimic the vocal identity of a target speaker. Modern deepfake tools can operate in real time, increasing the feasibility of live impersonation (Ren et al., 2019, 2020; Kim et al., 2020; Kaneko and Kameoka, 2018; Kameoka et al., 2018; Qian et al., 2019; Li et al., 2023; Qin et al., 2023; Wang et al., 2024). Recent advancements in zero-shot voice cloning enable attackers to generate convincing deepfakes with minimal reference audio, which lowers the barrier to entry. Currently, researchers produce CMs to prevent such attacks (Almutairi and Elgibreen, 2022).
One of the earliest demonstrations of AI-powered voice spoofing was conducted by Saxena et al. (Shonesy, 2015). Using a basic voice morphing tool, they were able to reconstruct a target speaker’s voice with only a few minutes of enrollment data. The spoofed voices were tested against both automated speaker verification systems and human listeners. Results revealed significant vulnerabilities: SOTA SRS failed to reject 80–90% of spoofed attempts, while human listeners correctly detected fake voices only about 50% of the time—particularly when the attacker mimicked familiar individuals.
Building on this, Alali et al. (Alali and Theodorakopoulos, 2025) explored partial voice attacks in real-world contexts. Their study demonstrated that both humans and SRS remain vulnerable when exposed to partial speech generated through zero-shot TTS and VC models. The findings highlighted a high success rate in bypassing verification mechanisms and deceiving listeners. However, the study’s broad evaluation across different attacks and platforms lacked a deeper analysis of specific conditions, such as noisy environments, language variations, or system-level defenses, limiting its applicability for improving individual systems.
@@ 2.5.1. White-box Attacks
Early work by Liu et al. (Liu et al., 2019) applied white-box adversarial techniques, notably the Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD), to CM systems. Even minimal perturbations (with $\epsilon$ below human perceptibility) caused error rates to spike dramatically, reaching 85–95% in some cases. These findings showed that, when attackers have access to model gradients, the CM’s ability to detect spoofed speech can be almost completely undermined, effectively reducing performance to random guessing.
Expanding on this, Gomez-Alanis et al. (Gomez-Alanis et al., 2022a) introduced the Adversarial Biometrics Transformation Network (ABTN), a trainable generator that optimizes perturbations using both CM and ASV losses. Compared to FGSM, ABTN produced perturbations that generalized better across spoof types while still preserving speaker identity, thereby fooling CMs without triggering alarms in ASV systems. The method improved attack equal error rates (EER) by 27–51% (absolute) across ASVspoof 2019 logical and physical access protocols (et al., 2019).
In a related study, Alanis et al. (Gomez-Alanis et al., 2022b) leveraged a generative adversarial network (GAN) to deceive both CM and speaker verification (SV) systems while maintaining speaker identity. Although highly effective, this approach required access to the target or a surrogate model and concentrated solely on attack generation, without proposing defensive strategies—limiting its practical applicability.
@@ 2.5.2. Black-box Attacks
In the domain of black-box adversarial attacks, Hai et al. (Hai et al., 2023) proposed SiFDetectCracker, a framework targeting CMs. Their method generates perturbations for speaker-irrelevant features (SiFs), specifically the background noise and the mute parts of an audio sample. The framework was evaluated against several advanced detectors, including RawNet2(Tak et al., 2021b), RawGAT-ST(Tak et al., 2021a), RawPC-Darts (Ge et al., 2021), and Deep4SNet (Ballesteros et al., 2021), using the ASVspoof 2019 dataset (et al., 2019). The attack proved highly effective, achieving an average success rate of 82.2% by showing that many detection systems are sensitive to these non-speech features. While the paper presents a novel attack vector, ablation studies showed that the attack’s performance drops significantly when either time or noise perturbations are removed, indicating a reliance on both factors. Furthermore, the transferability between models with different architectures was low, a limitation our work aims to address.
While the work by Kassis and Hengartner (Kassis and Hengartner, 2023) presents a novel methodology for bypassing VAS, its real-world applicability is limited by key limitations in its evaluation. Their results against combination systems rely on a cumulative metric, which is not descriptive enough to show the state or the actual effectiveness of their attack against these more complex defenses. Furthermore, the attack’s performance dropped significantly when tested against SOTA anti-spoofing systems, like RawPC-Darts (Ge et al., 2021) and RawGAT-ST (Tak et al., 2021a), raising doubts about its effectiveness against modern countermeasures. Finally, the designated-app attack vector relies on the significant assumption that an attacker can use a rooted phone to inject audio, a prerequisite that is often blocked by the security measures of financial applications.
@@ 3.1. Threat Model
  • Sample Generation: They can generate a spoofed audio sample using any voice-cloning tool (e.g., Fish Speech APIs (Liao et al., 2024) or methods used in ASVspoof 2019 (et al., 2019)).
@@
  • Score and Label: The system returns both a continuous confidence score and the final decision (like Azure SV (Microsoft, )).
@@ 3.2. Overview
As shown in Figure 2, the attack pipeline starts by generating a base spoofed audio. This is accomplished by feeding a victim’s voice sample and a target prompt into a voice synthesis model. Our framework is tool-independent—the attacker can employ Fish Speech APIs (Liao et al., 2024), ASVSpoof 2019 challenge tools (et al., 2019), or any modern voice cloning system.
@@ 4. Experimental Setup
Our experimental setup leverages the LibriSpeech (Panayotov et al., 2015) and ASVSpoof2019 (et al., 2019) datasets to assess system performance. We benchmark our proposed attacks against three state-of-the-art anti-spoofing models: RawNet2 (Tak et al., 2021b), RawGAT-ST(Tak et al., 2021a), and RawPC-Darts(Ge et al., 2021). For the voice authentication stage, we employ three distinct systems: X-Vectors (Snyder et al., 2018), DeepSpeaker (Li et al., 2017), and the commercial Azure Speaker Verification API (Microsoft, ). All experiments were conducted on a single NVIDIA T4 GPU. The evaluation is structured to analyze performance under various configurations: